# Resolving UnderEdit & OverEdit with Iterative & Neighbor-Assisted Model Editing

### v1

@@ B.1 Iterative model editing
Currently, our implementation requires running an additional iteration to compute $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ for iteration $k$. As a result, $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ for the 5th iteration is not reported in Tables 6, 7, 8, and 9. We are updating our code to compute $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ at the end of the $k$th iteration without requiring the next iteration. This update involves running only the Optimization stage with a single gradient step, using the initialization vector, before proceeding to the next iteration.
@@ B.2 Neighbor-assisted model editing
This approach also introduced an additional constraint on the edit data points: each data point must have at least two neighboring knowledge samples—one for editing and the others for evaluation. Unfortunately, the ZsRE dataset contains only one neighboring knowledge sample per data point, restricting us to the COUNTERFACT dataset. Even within COUNTERFACT, only a limited number of samples met the required constraints. The number of qualifying samples varied depending on the model, as shown in Table 2.
Appendix C Hardware Details
@@
Appendix D Iterative model editing results
@@
Appendix E Neighbor-assisted model editing results
@@
Our current implementation of neighbor-assisted model editing does not use prefixes. To analyze its behavior with prefixes, we conducted an additional set of experiments. The results, presented in Table 9, compare prefix-based neighbor-assisted model editing with its prefix-free counterpart and iterative model editing. A detailed analysis is provided in Section 6.2.

### v2

@@ B.1 Iterative model editing
Currently, our implementation requires running an additional iteration to compute $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ for iteration $k$. As a result, $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ for the 5th iteration is not reported in Tables 9, 10, 11, and 12. We are updating our code to compute $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$ at the end of the $k$th iteration without requiring the next iteration. This update involves running only the Optimization stage with a single gradient step, using the initialization vector, before proceeding to the next iteration.
@@ B.2 Neighbor-assisted model editing
This approach also introduced an additional constraint on the edit data points: each data point must have at least two neighboring knowledge samples—one for editing and the others for evaluation. Unfortunately, the ZsRE dataset contains only one neighboring knowledge sample per data point, restricting us to the COUNTERFACT dataset. Even within COUNTERFACT, only a limited number of samples met the required constraints. The number of qualifying samples varied depending on the model, as shown in Table 3. In neighbor-assisted model editing, exactly one neighboring example was used for each edit. Additionally, an average of 2, 4, and 4 neighboring examples (as reported in the Table 5) were used for evaluating each edit. Generating multiple neighbor knowledge using ChatGPT (Meng et al., 2023a) is straightforward and convenient. However, scaling this to all possible neighbors for a single piece of knowledge is a highly challenging research problem that falls beyond the scope of our paper.
Model Average evaluation neighboring samples (ceil)
GPT-2 XL 2
GPT-J 4
Llama-2-7B 4
Table 5: Average neighboring sample used to evaluate each edit in neighbor-assisted editing
@@
Appendix C Hardware Details
k Efficacy ($\uparrow$) Generalization ($\uparrow$) Specificity ($\uparrow$) Score ($\uparrow$) Perplexity ($\downarrow$)
Accuracy Success Accuracy Success Accuracy Success Accuracy Success ME-PPL-50 $p(\theta_{k},\bar{h}_{t,k}^{l_{c}})$ $p(\theta_{k+1},\hat{h}^{l_{c}}_{t,k})$
1 79.00% 93.00% 22.00% 65.00% 75.00% 99.00% 42.00% 83.00% 57.98 1.02 10816.41
2 89.00% 98.00% 35.00% 76.00% 61.00% 98.00% 53.00% 89.00% 84.28 1.02 3700.56
3 68.00% 95.00% 28.00% 72.00% 54.00% 97.00% 44.00% 86.00% 174.62 1.02 64858.83
4 41.00% 85.00% 19.00% 68.00% 47.00% 97.00% 31.00% 81.00% 348.5 1.02 2254146.29
5 20.00% 76.00% 11.00% 64.00% 37.00% 97.00% 17.00% 76.00% 602.14 1.02 5827805.82
Table 7: Iterative Spread model editing results on COUNTERFACT with MEMIT on GPT-2 XL for 5 iterations. Unlike Spread, Optimization was run only in first iteration.
Appendix D Iterative Spread
In Figure 3, we observe little to no change in the perplexity of $\bar{h}^{l_{c}}_{t,k}$ across iterations. However, we argue that despite similar perplexity values, the actual representations of $\bar{h}^{l_{c}}_{t,k}$ differ at each step. This is because $\theta_{k+1}$ is computed based on $\bar{h}^{l_{c}}_{t,k}$, which itself is derived from $\theta_{k}$, as described in Section 3.1. Due to this recursive dependency, every update to $\theta_{k}$ induces a corresponding change in $\bar{h}^{l_{c}}_{t,k}$. Since the overall editing goal remains consistent across iterations, the magnitude of change required in $\theta_{k}$ naturally diminishes over time, resulting in a sequence of distinct yet converging hidden states.
To further explore this, consider a scenario where $\bar{h}^{l_{c}}_{t,k}$ remains constant across iterations. In such a case, one might skip repeated calls to Optimization and only perform Spread, potentially saving computation without sacrificing performance. However, we hypothesize that this would lead to overfitting due to repeated use of the same update. To test this, we conducted an additional experiment using iterative Spread with MEMIT on GPT-2 XL (COUNTERFACT dataset), where Optimization was run only in the first iteration. The results, shown in Table 7, reveal performance degradation and an increase in ME-PPL-50. Although we observe a temporary performance improvement in iteration 2, likely due to overfitting, it is followed by continued degradation, and the editing process fails to meet its stopping criterion. This underscores the importance of running Optimization in every iteration and indicates that it generates distinct target hidden states at each step.
Appendix E Factors Contributing Model Collapse
Existing literature attributes model collapse during editing to two main factors: (1) the sequential nature of edits (Gupta et al., 2024b; Yang et al., 2024), and (2) characteristics of the specific examples being edited (Yang et al., 2024). Since iterative model editing falls under the first category, one might expect it to be susceptible. However, prior work shows that collapse typically occurs after thousands of edits, e.g., up to 3k in AlphaEdit Fang et al. (2025) and 10k in ENCORE Gupta et al. (2025), whereas our iterative approach involves only a few single-digit iterations. This makes collapse due to iteration count alone unlikely.
Given that all models in our study were evaluated on the same dataset, we can reasonably rule out the second factor. Furthermore, since no GPT models, and not even LLaMA-3.1 (8B), exhibited similar collapse behavior, we hypothesize a third contributing factor: model-specific characteristics that may stem from differences in training strategies. Finally, the editing algorithm itself may serve as a fourth contributing factor. For example, MEMIT’s less targeted weight updates (compared to PMET) may introduce excessive parameter shifts, as discussed in Appendix A.
In summary, while iterative model editing is unlikely to cause collapse on its own, interactions with model-specific or algorithm-specific factors may trigger instability. Although investigating collapse is not the primary focus of this work, our findings offer a useful foundation for future research in this area.
Appendix F Additional Analysis
This section dives deeper into the proposed methods. Specifically, we aim to understand how iterative model editing addresses UnderEdit, the impact of prefixing in neighbor-assisted model editing for resolving OverEdit, and the effectiveness of the stopping criteria.
F.1 How does iterative model editing address UnderEdit?
We hypothesize that the iterative model editing approach can reduce the number of UnderEdit cases. To test this hypothesis, we identified UnderEdit examples in GPT-2 XL edited with MEMIT on COUNTERFACT after the first iteration, i.e, edits $(s_{i},r_{i},o_{i},o^{*}_{i})$ where $f_{\theta_{2}}(x_{i})[o_{i}^{*}]<f_{\theta_{2}}(x_{i})[o_{i}]$. We then tracked $|\Delta p_{k}|$ and efficacy accuracy across subsequent iterations. Figure 4 illustrates the results of the analysis. We observe that the rate of $|\Delta p_{k}|$ decrease and accuracy improving over iterations is much more pronounced for the UnderEdit examples. This observation tells us that multiple iterations of Spread is the larger contributors to getting higher performance.
Accuracy Accuracy Accuracy Accuracy
Unedited 0 1.00% 1.00%
MEMIT 4 99.00% 38.00% 52.00% 54.00%
NA_MEMIT 4 99.00% 36.00% 86.00% 60.00%
NAP_MEMIT 4 99.00% 37.00% 84.00% 61.00%
PMET 8 99.00% 31.00% 67.00% 52.00%
NA_PMET 9 98.00% 29.00% 85.00% 53.00%
NAP_PMET 8 98.00% 30.00% 84.00% 54.00%
F.2 How prefixes influence neighbor-assisted model editing behavior?
Using random prefixes aid generalization across contexts in the memory editing process when only the target knowledge edit is known. So we pose the question, does adding random prefixes to the neighbor knowledge prompts help prevent OverEdit? To answer this question we run an experiment by adding random prefixes to the neighboring knowledge used during edit. Table 8 shows an increase in specificity accuracy [^6] The full results with success on each measurement and perplexity performance are in Table 12. However, this increase is less compared to the improvement of using neighbor-assisted editing versus no neighbor-assist. Regardless, the prefixed neighbor-assisted edits (NAP_) achieved better overall performance, denoted by Accuracy in Score, due to a boost in generalization.
Appendix G Iterative model editing results
@@
Appendix H Neighbor-assisted model editing results
@@
Our current implementation of neighbor-assisted model editing does not use prefixes. To analyze its behavior with prefixes, we conducted an additional set of experiments. The results, presented in Table 12, compare prefix-based neighbor-assisted model editing with its prefix-free counterpart and iterative model editing. A detailed analysis is provided in Section F.2.